# Predictive Data Selection: The Data That Predicts Is the Data That Teaches

### v2

@@ 1 Introduction
Our approach shares a similar intuition with (Thrush et al., 2024), while we are motivated independently by Huang et al. (2024). However, this work differs from Thrush et al. (2024) in several key aspects: (1) While they compute perplexity correlations for data domains (e.g., wikipedia.com) and select domains, our method operates at a much finer granularity, directly operating on and selecting individual documents. As we will demonstrate in §3.4 and §4.3, this finer granularity leads to significant differences in the selected data and, consequently, in the final results; (2) They performed very limited experiments with a 160M-size model and compared with several baselines on limited benchmarks, while we train models with up to 3B parameters on up to 100B tokens and evaluate on 17 diverse benchmarks, a more realistic scale in a typical pretraining setting.
@@ 2.2 Data’s Predictive Strength
Beyond considerations of simplicity, we argue that this ranking-based correlation is more robust to noise compared to the Pearson correlation used by Huang et al. (2024). Loss computation on a single document can be sensitive to noisy or anomalous text, especially when the document is short. In such cases, numerical Pearson correlation, which reacts sharply to minor variations in loss values, may not be a suitable metric. This instability can, to some extent, be mitigated by grouping documents first and selecting based on these groups, as practiced by Thrush et al. (2024), who group and select data by domain URLs. However, such a grouping step introduces additional human heuristics into the process that we think may be suboptimal, as we will verify empirically in §3.4.
@@ 2.3 The Overall Framework
Our core hypothesis is that the data on which the losses can help predict the performance well is the data that can contribute to training effectively, thus the documents’ predictive strength scores as defined in Eq. 1 are our metric to select data. However, computing Eq. 1 for every document is expensive since it requires computing losses of several models on all the data, thus we follow Li et al. (2024) and only select a small seed set of documents that are later used to train a fastText classifier (Joulin et al., 2017). As shown in Figure 2, we first randomly sample a small subset of data from the pre-training corpus, which is kept separate from the data used for later training, to compute the normalized losses. Specifically, we pick the 3,000 most frequent domains (e.g., wikipedia.org) in the corpus and randomly sample 300 examples for each domain for a wide coverage over the pre-training corpora, thus obtaining 900K samples in total. Then we choose the models from the Llama 1 and 2 series  (Touvron et al., 2023a, b) ranging from 7B to 65B parameters – in total 6 models are used to compute the predictive strength for the data. We run inference of these models on the 900K documents to obtain the loss ranks of these models on each document respectively. Next, we obtain the downstream score ranks by directly using the averaged scores on a diverse set of 12 benchmarks from Huang et al. (2024).
@@ 3.2 Baselines
To make a fair comparison, we keep the data pool and training setting of baselines the same as PreSelect, such as model architecture and training hyper-parameters, while only the data selection part is different. We consider the following baselines: (1) Random, which randomly selects the documents; (2) PPL filtering, which keeps low-perplexity documents following CCNet (Wenzek et al., 2020); (3) FineWeb-Edu (Penedo et al., 2024b), which scores the education level of documents with an LLM and then train a small scorer to identify and select educational contents (Penedo et al., 2024b); (4) PPL Correlation (Thrush et al., 2024), which utilizes perplexity correlation to calculate domain-level correlation and selects domains. As discussed before in §1, this baseline is the most relevant to PreSelect while our approach does not rely on human heuristics to group documents and operates at the document level directly; and (4) DCLM (Li et al., 2024), which is the state-of-the-art pretraining data selection method that trains the fastText scorer using supervised fine-tuning (SFT) data as the positive data. For FineWeb-Edu and DCLM, we directly use their released BERT and fastText scorer to select documents in our pool. Due to the high cost of running all the baselines across all settings, we mainly include all of them in our setting of 1B model training for 30B tokens. In other experiment scales, we only compare to DCLM which is superior to other baselines. We provide more details about the model architecture and training hyper-parameters in Appendix B.3 and Appendix B.4 respectively.
@@ 3.4 Main Results
Compared with Perplexity Correlation which selects domains based on the correlation of each domain, it only achieves relatively marginal improvements over random selection and is significantly underperformed PreSelect by 4.8%. This is because by operating at the domain level, it may inadvertently include low-quality data from the high-correlation domains while overlooking high-quality data from unselected domains. As we will further discuss in § 4.3, the removal of many domains leads to a significantly decreased diversity, thus, some tasks such as Arc-Easy would even have a significant drop.
@@
PPL Correlation www.theguardian.com (1.12), issuu.com (1.11), www.nifty.org (0.90), www.fanfiction.net (0.83), openjurist.org (0.68), www.beeradvocate.com (0.68), www.nytimes.com (0.65), www.literotica.com (0.58), www.washingtonpost.com (0.54), www.slideshare.net (0.54), www.mdpi.com (0.48), archive.org (0.45), www.businessinsider.com (0.43), www.barnesandnoble.com (0.43), www.forbes.com (0.43)
@@ 4.2 FastText Feature Contribution
As shown in Table 3, for DCLM, the baseline that also uses a fastText classifier, the uni-gram features with high influence scores include (Why, 672), (Answer:, 620), (assistant., 567), (Question:, 448), (Describe, 368)… It is clearly observed that most positive uni-gram features have supervised fine-tuning data patterns and if the data in the pretraining corpus has such uni-gram, they have a high probability to be selected. While for PreSelect, the uni-gram features with high influence scores include (^ , 338), (MIT, 241), (Math, 194), (answer, 188), which widely covers many features appeared in code, education, and question answering domains. For example, “^” exists in many code snippets, “MIT” and “Math” exist in many education-related examples. The full learned positive and negative uni-grams alongside their influence score can be found in appendix A.3.2 which provides more insights into what word patterns are preferred by different methods.
@@ 4.3 Distribution of Selected Data
FineWeb-Edu aims to select education-related contents which prioritize many education-related domains such as wikisource.org, britannica.com, phys.org, ncbi.nlm.nih.gov, while down-sampled those general utility domains. This makes FineWeb-Edu achieve great performance on examination-related benchmarks such as ARC, hurting the performance on understanding such as LAMBADA. DCLM aims to select SFT-like data and is clearly reflected in the selected data, where a large number of question-answer related domains are up-sampled such as physicsforums.com, stackexchange.com, reddit.com, stackoverflow.com. These SFT-like data brings DCLM significant advantages over random-sampled baseline on many downstream tasks. Perplexity correlation selects domains based on domain-level correlation, from the highest correlation domains until filling up token budgets. This means it will discard all the remaining domains with relatively medium/low correlations once sufficient training tokens are obtained (e.g. only  1500 domains are selected to fill up the required training tokens). As shown in Table 4 row 4, many high-quality domains such as wikipedia.org may be dropped according to medium correlation and automatically upsampled each single selected domain, such as theguardian.com weights from 0.12% in original data distribution to 1.12% in perplexity correlation. This significantly hurt the data diversity and quality, even leading to a performance drop on HellaSwag, Arc-Easy compared to random selection, as shown in Appendix Table 11.

### v3

@@ 1 Introduction
While we are motivated independently by Huang et al. (2024), our approach shares similar intuition with Perplexity Correlation (Thrush et al., 2024) and can be viewed under the framework of Perplexity Correlation, which is the first to explore domain-level correlation-based methods in data selection. However, this work differs from Thrush et al. (2024) in several key aspects: (1) While they compute perplexity correlations for data domains (e.g., wikipedia.com), our method operates at a much finer granularity, directly operating on and selecting individual documents. [^2] We note that Thrush et al. (2024) preregistered document-level experiments as future tasks, which were not present as we released the first version of this paper. The updated version of Perplexity Correlation added document-level results concurrent to this work, but the final results are very different as we will dicuss in §2.3. As we will demonstrate in §3.4 and §4.3, this finer granularity leads to significant differences in the selected data and, consequently, in the final results; (2) Perplexity correlation suggested using a large number of (e.g., around 90) open-sourced models to approximate the correlation. However, we argue that using many open-sourced LLMs brings potential evaluation noises when dealing with different families of the models and causes computation inefficiency. Thus in our design, we employ 6 Llama models only during data selection. (3) They performed very limited experiments with a 160M-size model and compared with several baselines on limited benchmarks, while we train models with up to 3B parameters on up to 100B tokens and evaluate on 17 diverse benchmarks, a more realistic scale in a typical pretraining setting. The updated, concurrent version of Thrush et al. (2024) further added 1B-scale experiments assesed on comprehensive benchmarks, where the results indicate good performance on raw data pools but no improvement on high-quality, well-processes data pools – an observation distinct from our empirical results. PreSelect targets a well de-duplicated, high-quality pre-training corpus – which we believe represents a realistic setting – and shows the effectiveness of selecting data from high-quality data pools in §3.4.
@@ 2.2 Data’s Predictive Strength
Beyond considerations of simplicity, we argue that this ranking-based correlation is more robust to noise compared to the Pearson correlation used by Huang et al. (2024) in document-level calculation. As in prior, Perplexity Correlation (Thrush et al., 2024) has explored ranking-based metric to estimate correlation coefficient in domain-level. Loss computation on a single document can be sensitive to noisy or anomalous text, especially when the document is short. In such cases, numerical correlation estimation such as Pearson correlation, which reacts sharply to minor variations in loss values, may not be a suitable metric. This instability can, to some extent, be mitigated by grouping documents first and selecting based on these groups. However, such a grouping step introduces additional human heuristics into the process that we think may be suboptimal, as we will verify empirically in §3.4.
@@ 2.3 The Overall Framework
Our core hypothesis is that the data on which the losses can help predict the performance well is the data that can contribute to training effectively, thus the documents’ predictive strength scores as defined in Eq. 1 are our metric to select data. However, computing Eq. 1 for every document is expensive since it requires computing losses of several models on all the data, thus we follow Li et al. (2024) and only select a small seed set of documents that are later used to train a fastText classifier (Joulin et al., 2017). As shown in Figure 2, we first randomly sample a small subset of data from the pre-training corpus, which is kept separate from the data used for later training, to compute the normalized losses. Specifically, we pick the 3,000 most frequent domains (e.g., wikipedia.org) in the corpus and randomly sample 300 examples for each domain for a wide coverage over the pre-training corpora, thus obtaining 900K samples in total. Then we choose the models from the Llama 1 and 2 series  (Touvron et al., 2023a, b) ranging from 7B to 65B parameters – in total 6 models are used to compute the predictive strength for the data. We do not choose various models from different families to compute the predictive strength on purpose – in our preliminary experiments, we found that models from various families introduce significant evaluation noises, for example, they are sensitive to prompts and the prompt that is good for one model may not be suitable for another. This issue is particularly salient given that we are evaluating base models that are generally more sensitive to prompts than instruction-tuned models. Such noises would dramatically affect the computation of predictive strength, and thus we found that data selected this way could not outperform a random data selection baseline in our initial trials. This design differs from Thrush et al. (2024) which employs diverse models to estimate perplexity correlation. Such distinction probably explains why perplexity correlation fails to improve random data selection baselines significantly on a high-quality, pre-filtered data pool, as shown in the concurrent version of their paper. We run inference of these models on the 900K documents to obtain the loss ranks of these models on each document respectively. Next, we obtain the downstream score ranks by directly using the averaged scores on a diverse set of 12 benchmarks from Huang et al. (2024).
@@ 3.2 Baselines
To make a fair comparison, we keep the data pool and training setting of baselines the same as PreSelect, such as model architecture and training hyper-parameters, while only the data selection part is different. We consider the following baselines: (1) Random, which randomly selects the documents; (2) PPL filtering, which keeps low-perplexity documents following CCNet (Wenzek et al., 2020); (3) FineWeb-Edu (Penedo et al., 2024b), which scores the education level of documents with an LLM and then train a small scorer to identify and select educational contents (Penedo et al., 2024b); (4) PPL Correlation (DD) (Thrush et al., 2024), which utilizes domain-level correlation estimation and performs domain-level selection. (5) PPL Correlation (DP) (Thrush et al., 2024), which utilizes domain-level correlation estimation and trains a fasttext based on domain-level positive/negative samples to perform page-level selection. As discussed before in §1, these two baselines are the most relevant to PreSelect while our approach does not rely on human heuristics to group documents and operates at the document level directly; and (6) DCLM (Li et al., 2024), which is the state-of-the-art pretraining data selection method that trains the fastText scorer using supervised fine-tuning (SFT) data as the positive data. For FineWeb-Edu and DCLM, we directly use their released BERT and fastText scorer to select documents in our pool. Due to the high cost of running all the baselines across all settings, we mainly include all of them in our setting of 1B model training for 30B tokens. In other experiment scales, we only compare to DCLM which is superior to other baselines. We provide more details about the model architecture and training hyper-parameters in Appendix B.3 and Appendix B.4 respectively.
@@ 3.3 Evaluation Settings
PPL Correlation (DP)(pos) (Archives, 347.5), (Navigation, 299.7), (Commenting, 288.5), (Section, 280.3), (E-mail, 276.1), (Played:, 263.9), (HB, 249.7), (Lyrics, 240.4), (Received, 231.4), (Gamercard, 226.7), (+$0$, 226.6), (Discussion:, 224.3), (Sleeps, 219.0), (Playing:, 216.9), (Forgot, 216.3)
PPL Correlation (DP)(pos) (Login, -459.9), ((permalink), -367.6), ((IANS), -358.0), (Images), -344.0), (macrumors, -343.3), ((credit:, -320.8), (Shipping, -314.9), (Subscription, -293.5), (Defines, -291.0), (Number:, -286.5), (Lainey, -279.2), (Maine, -278.0), (Tutors, -277.6), (Nearby, -269.7), (By:, -269.5)
@@ 3.4 Main Results
Compared with Perplexity Correlation (DD)which selects domains based on the correlation of each domain, it only achieves relatively marginal improvements over random selection and is significantly underperformed PreSelect by 4.8%. As we will further discuss in § 4.3, the removal of many domains leads to a significantly decreased diversity, thus, some tasks such as HellaSwag would even have a significant drop. In addition, we compare with Perplexity Correlation (DP) which perform page-level selection using fastText trained with domain-level positive/negative samples. We see that though it has 1.2% improvements over random baseline, it still underperformed PreSelect by 4.1%, showing the effectivness of using example-level correlation estimation. This difference is because by operating at the domain level, it may inadvertently include low-quality data from the high-correlation domains while overlooking high-quality data from unselected domains when building the fastText classifier.
@@
PPL Correlation (DD) www.theguardian.com (1.12), issuu.com (1.11), www.nifty.org (0.90), www.fanfiction.net (0.83), openjurist.org (0.68), www.beeradvocate.com (0.68), www.nytimes.com (0.65), www.literotica.com (0.58), www.washingtonpost.com (0.54), www.slideshare.net (0.54), www.mdpi.com (0.48), archive.org (0.45), www.businessinsider.com (0.43), www.barnesandnoble.com (0.43), www.forbes.com (0.43)
PPL Correlation (DP) (www.fanfiction.net, 0.36), (www.opentable.com, 0.32), (www.ncbi.nlm.nih.gov, 0.30), (www.alldiscountbooks.net, 0.29), (www.hindawi.com, 0.26), (pubmedcentralcanada.ca, 0.26), (www.rockpapershotgun.com, 0.23), (www.literotica.com, 0.22), (link.springer.com, 0.20), (politicalticker.blogs.cnn.com, 0.20), (sixstat.com, 0.19), (openjurist.org, 0.19), (www.homeaway.com, 0.17), (www.nifty.org, 0.17), (clinicaltrials.gov, 0.16)
@@ 4.2 FastText Feature Contribution
As shown in Table 3, for DCLM, the baseline that also uses a fastText classifier, the uni-gram features with high influence scores include (Why, 672), (Answer:, 620), (assistant., 567), (Question:, 448), (Describe, 368)… It is clearly observed that most positive uni-gram features have supervised fine-tuning data patterns and if the data in the pretraining corpus has such uni-gram, they have a high probability to be selected. While for PreSelect, the uni-gram features with high influence scores include (^ , 338), (MIT, 241), (Math, 194), (answer, 188), which widely covers many features appeared in code, education, and question answering domains. For example, “^” exists in many code snippets, “MIT” and “Math” exist in many education-related examples. These representative uni-gram features does not show in domain-level correlation estimation method, Perplexity Correlation (DP), where various diverse meaningless features are learned. We attribute this to the noise in the domain-level correlation calculation, where inadvertently include low-quality data from the high-correlation domains while overlooking high-quality data from unselected domains. The full learned positive and negative uni-grams alongside their influence score can be found in appendix A.3.2 which provides more insights into what word patterns are preferred by different methods.
@@ 4.3 Distribution of Selected Data
FineWeb-Edu aims to select education-related contents which prioritize many education-related domains such as wikisource.org, britannica.com, phys.org, ncbi.nlm.nih.gov, while down-sampled those general utility domains. This makes FineWeb-Edu achieve great performance on examination-related benchmarks such as ARC, hurting the performance on understanding such as LAMBADA. DCLM aims to select SFT-like data and is clearly reflected in the selected data, where a large number of question-answer related domains are up-sampled such as physicsforums.com, stackexchange.com, reddit.com, stackoverflow.com. These SFT-like data brings DCLM significant advantages over random-sampled baseline on many downstream tasks. Perplexity correlation (DD) selects domains based on domain-level correlation, from the highest correlation domains until filling up token budgets. This means it will discard all the remaining domains with relatively medium/low correlations once sufficient training tokens are obtained (e.g. only  1500 domains are selected to fill up the required training tokens). As shown in Table 4 row 4, many high-quality domains such as wikipedia.org may be dropped according to medium correlation and automatically upsampled each single selected domain, such as theguardian.com weights from 0.12% in original data distribution to 1.12% in perplexity correlation (DD). This significantly hurt the data diversity and quality, even leading to a performance drop on HellaSwag, Arc-Challenge compared to random selection, as shown in Appendix Table 11.
Perplexity correlation (DP) utilizes domain-level correlation estimation and train a fasttext based on domain-level positive/negative samples to perform page-level selection. This mitigate the problems in Perplexity correlation (DD) to some extent, but it suffers from the noise inside each domains as we discussed in §4.2. This also reflected in final selected data where the learned features fail to upsample high-quality domains as shown in Table 4.